### v3

# pyCFS-data: Data Processing Framework in Python for openCFS

Abstract
Many numerical simulation tools have been developed and are on the market, but there is still a strong need for appropriate tools capable of simulating multi-field problems, especially in aeroacoustics. Therefore, openCFS provides an open-source framework for implementing partial differential equations using the finite element method. Since 2000, the software has been developed continuously. The result is openCFS (before 2020, known as CFS++ Coupled Field Simulations written in C++). In this paper, we present pyCFS-data, a data processing framework written in Python to provide a flexible and easy-to-use toolbox to access and manipulate, pre- and postprocess data generated by or for usage with openCFS.
Keywords Data processing framework $\cdot$ Python $\cdot$ Open Source FEM Software $\cdot$ Multiphysics Simulation $\cdot$ openCFS
1 Introduction
Within this contribution, we concentrate on the pyCFS[5] module pyCFS-data, a data processing framework for openCFS [9]. It consists of three main components, which are organized in same-titled submodules
  • io
  • operators
  • extras
where the io submodule focuses on reading, writing, and accessing data in HDF5 file format as used in openCFS. The operators submodule contains functionality for data pre-, and post-processing based on data structures defined in io submodule. The extras submodule contains methods to interact with data stored in different file formats.
The module can be easily installed, including all dependencies, via the pyPI [^1] https://pypi.org/project/pyCFS/
pip install pycfs
  Detailed instructions for installation inside a virtual environment and developer setup can be found at https://opencfs.gitlab.io/pycfs/installation.html
When establishing an XML file for CFS-Data, it is fundamental that the pipeline, existing of different CFS-Data filters, is closed. The pipeline has to start with the step value definition and has to be followed by the input filter and end with the output filter. In between, multiple filters can be added, serial or parallel.
In this contribution, we discuss the capability of each submodule of pyCFS-data in more detail. Comprehensive and up-to-date documentation of the library can be found in the pyCFS API Documentation [^2] https://opencfs.gitlab.io/pycfs/generated/pyCFS.data.html.
2 I/O of openCFS based HDF5 files (io)
Module for input and output operations of openCFS type mesh and result data files. The simplest way of reading and writing files is via the top level utility functions read_file, read_mesh, read_data, write_file. A simple example reads
from pyCFS.data import io
mesh, results = io.read_file(’file.cfs’)
io.write_file(’file_out.cfs’, mesh, results)
, reading a file and writing the content to a new file. Figure 1 gives an overview of the I/O module’s structure.
Figure 1: Organization of classes for data structures.
2.1 Reader class (CFSReader)
Base class for all reading operations. A simple example of usage could look like the following
from pyCFS.data.io import CFSReader
with CFSReader(’file.cfs’) as f:
mesh = f.MeshData
results = f.MultiStepData
, which reads the file file.cfs into the data objects mesh and results, containing the description of the computational grid and all result data of the default analysis step (default is $1$), respectively.
2.2 Writer class (CFSWriter)
Base class for all writing operations. A simple example of usage could look like the following
from pyCFS.data.io import CFSReader, CFSWriter
with CFSReader(’file_in.cfs’) as f:
mesh = f.MeshData
results = f.MultiStepData
with CFSWriter(’file_out.cfs’) as f:
f.create_file(mesh=mesh, result=results)
, which reads the file file_in.cfs into the data objects mesh and results, containing the description of the computational grid and result data, respectively. Thereafter, this information is written to a newly created file file_out.cfs.
2.3 Data structures (CFSMeshData,CFSResultContainer, etc.)
The structure of data is organized as depicted in Fig. 1.
CFSMeshData: Data structure containing mesh definition. Its attributes and properties are depicted in Fig. 2.
Figure 2: Structure of CFSMeshData class. Attributes are highlighted in orange, Properties in purple, and respective data types in grey.
CFSMeshInfo: Data structure containing mesh information.
CFSRegData: Data structure containing mesh region definition. Its attributes and properties are depicted in Fig. 3.
Figure 3: Structure of CFSRegData class. Attributes are highlighted in orange, Properties in purple, and respective data types in grey.
CFSResultContainer: Data structure containing result data (one object currently supports a single MultiStep only!). Its attributes and properties are depicted in Fig. 4.
Figure 4: Structure of CFSResultContainer class. Attributes are highlighted in orange, Properties in purple, and respective data types in grey.
CFSResultArray: Overload/Subclass of numpy.ndarray  [^3] https://numpy.org/doc/stable/reference/generated/numpy.ndarray.html adding attributes according to CFS HDF5 result data structure. Its additional attributes and properties are depicted in Fig. 5.
  Field data can be defined on the nodes or the cell centroids of a computational grid, while history data can also be defined on Region (ResType attribute), defining the required shape of the array data. Field data requires $(N,M,D)$ while history data requires $(N,D)$, with the number of steps $N$, the number of degrees of freedom (DOF) $M$ and the number of dimensions $D$. Data can be defined in the time or the frequency domain, based on the AnalysisType attribute.
  Dimension names can be defined optionally (DimNames attribute). If not specified, they are chosen automatically based on the shape of the data array.
  The attribute IsComplex can be defined optionally to ensure writing complex-valued data or enforce real-valued data. If not specified, it will be chosen automatically based on the corresponding analysis type (AnalysisType attribute).
Figure 5: Structure of CFSResultArray class. Attributes are highlighted in orange, Properties in purple, and respective data types in grey.
  The Quantity attribute can be chosen as any string in general. However, if the exported data will be the input of a subsequent openCFS simulation, openCFS variable names must be used for the declaration of field quantities. Thus, for the acoustic PDE, one of the following names must be chosen. General acoustic and fluid mechanic quantities: acouPressure acouVelocity acouPotential acoutIntensity fluidMechVelocity meanFluidMechVelocity fluidMechPressure fluidMechDensity fluidMechVorticity fluidMechGradPressure Aeroacoustic Source Terms: acouRhsLoad (general) acouRhsLoadP (Perturbed convective wave equation) vortexRhsLoad (Vortex sound theory) acouDivLighthillTensor (Lighthill’s acoustic analogy)
CFSResultInfo: Data structure containing result information for one result. Its attributes and properties are depicted in Fig. 6.
  The CFSResultInfo structure is redundant to the metadata contained within a CFSResultArray object!
Figure 6: Structure of CFSResultInfo class. Attributes are highlighted in orange, Properties in purple, and respective data types in grey.
3 Operators for data processing (operators)
Library of modules to perform various operations on objects in pyCFS-data data structures (see Sec. 2.3).
3.1 Interpolation operations (interpolators)
Module containing methods for interpolation operations.
3.1.1 Node2Cell
The node-to-cell interpolation operator takes nodal loads and connects them to the cell center, of the cell defined by those nodes
$e_{\square}=\frac{1}{n}\sum_{i=1}^{n}v_{i}\,.$ (1)
Thereby, $e_{\square}$ is the data assigned to the element, $\mathrm{n}$ is the number of nodes of the element, and $v_{i}$ are the nodal loads.
The following example shows this methodology by considering one linear triangular element:
Figure 7: Node to cell interpolator.
3.1.2 Cell2Node
The cell-to-node interpolation operator takes element loads and divides them onto the nodes that build the cell.
$v_{\square}=\frac{1}{n}\sum_{i=1}^{n}e_{i}\,.$ (2)
Thereby, $v_{\square}$ is the data located to the node, $n$ the number of adjacent elements to the node, and $e_{i}$ the data of cell $i$.
The following example shows this methodology by considering one linear triangular element:
Figure 8: Cell to node interpolator.
3.1.3 Nearest Neighbour
The implemented nearest neighbor interpolation operator makes use of inverse distance weighting (Shepard’s method).
Figure 9: Nearest neighbor interpolator using ’forward’ formulation.
Based on the defined number of neighbors $n$ from the source mesh, the nearest neighbors are searched, and their distance $r$ to the considered node is computed. Based on this distances the weights $w_{i}$ are computed as
$w_{i}=\left(\frac{R_{\mathrm{max}}-r_{i}}{R_{\mathrm{max}}r_{i}}\right)^{p}\,$, (3)
with $R_{\mathrm{max}}=1.01r_{\mathrm{max}}$ as being 1.01 times of the maximal distance $r_{\mathrm{max}}$, and $p$ as the interpolation exponent. Shepard stated $1\leq p\leq 3$. Increasing $p$ means that values that are further away are taken into account more. Finally, each value of each node $v_{1}$ that is taken into account is weighted to compute the new value $v_{\mathrm{n}}$ of the considered node
$v_{\mathrm{n}}=\sum_{i=1}^{n}\frac{w_{i}v_{i}}{\sum_{i=1}^{n}w_{i}}\,.$ (4)
Two search directions for the nearest neighbor search are implemented:
  • ’forward’: Nearest neighbors are searched for each point on the (coarser) target grid. This leads to a checkerboard if the target grid is finer than the source grid.
  • ’backward’: Nearest neighbors are searched for each point on the (coarser) source grid. This leads to overprediction if the source grid is finer than the target grid.
  By default, the formulation is chosen automatically based on the number of source and target points.
3.1.4 Projection-based interpolation
This interpolation operator projects points of the target mesh onto the source mesh and evaluates the result based on linear FE basis functions. A comprehensive description of the algorithm can be found in [14].
Figure 10: Projection-based interpolation. Sketch of an arbitrary source (blue) and target mesh (red) with indicated point projection from the target to the source mesh. [14]
Parameters include
  • proj_direction: Direction vector used for projection. It can be specified as constant or individually for each node. By default, the node normal vector (based on averaged neighboring element normal vectors) is used.
  • max_distance: Maximum projection distance. Lower values speed up interpolation matrix build and prevent projecting onto far surfaces.
  • search_radius: Radius of elements considered as projection targets. It should be chosen at least to the maximum element size of the target grid.
3.1.5 Radial basis function interpolation and differentiation
This module provides functions to interpolate data and compute gradients using radial basis functions (RBF). Implemented kernel functions are: Gaussian RBF kernel, Multiquadratic RBF kernel, and Wendland C2 RBF kernel.
The implemented methods include interpolation and gradient computation of a field quantity using global or local RBF approximation. Parameters include
  • kernel: Type of RBF kernel to use. Available kernels include gaussian, multiquadratic, and wendland_c2.
  • epsilon: Shape parameter for the used RBF kernel function
  • smoothing: Smoothing parameter for the kernel matrix
for the global interpolation and additionally for the local interpolation
  • neighbors: Number of nearest neighbors to use for local interpolation
  • min_neighbors: Minimum number of neighbors to use based on a local neighborhood radius $r=f_{r}\frac{1}{N}\sum_{i=1}^{N}d_{i}$ with factor $f_{r}$, and the distance of the nearest neighbor distances $d_{i}$
  • radius_factor: Factor $f_{r}$ to determine the local neighborhood radius.
3.2 Modal analysis tools (modal_analysis)
Common metrics used in modal analysis including the Modal Assurance Criterion (MAC), the Modal Scale Factor (MSF), and the Modal Complexity Factor (MCF).
3.3 Synthetic Noise Generation and Radiation (SNGR) source computation (sngr)
Computation of synthetic flow acoustic sources based on the theory of [1] for the SNGR with additional description found in [13]. The implementation is described in detail in [16].
3.4 Transformation operations (transformation)
Module containing methods for mesh/data transformation operations.
3.4.1 Extrude/Revolve mesh(extrude_mesh_region, revolve_mesh_region)
Create a 3D/2D mesh by extruding/revolving a 2D/1D base mesh along a path.
3.4.2 Transform mesh data (transform_mesh_data)
Perform a rigid body transformation of a mesh and its corresponding vector data.
3.4.3 Fit geometry (fit_mesh)
This script fits a region to a target region using rotation and translation transformations based on minimizing the squared distance of all source nodes to the respective nearest neighbor on the target mesh. The transformation parameters are x,y,z displacement, and Euler angles.
3.5 Spatial derivatives (derivatives)
Radial basis function based evaluation of spatial derivatives (see section˜3.1.5).
3.6 Time signal processing
Module containing methods for time domain data processing operations.
3.6.1 Time derivative (timeproc.time_derivative)
Compute the time derivative of a result array using a noise-robust differentiator. The time derivative of the desired quantity $\partial q/\partial t$ is computed by a smooth noise-robust differentiator, which supresses high frequencies and is precise on low frequencies, according to [3]. For an efficient and robust calculation the order of the differentiator is set to $N=5$ and the time derivative is calculated by
$\frac{\partial q}{\partial t}\approx\frac{2(q_{1}-q_{-1})+q_{2}-q_{-2}}{8\Delta t}\,$, (5)
where the index of $q$ defines the time step relative to the time step of which the derivative is calculated and $\Delta t$ is the time step size.
Parameters include
  • boundary_treatment: The start and end of the time series can either be removed (remove), left untreated ((None)), or treated by one-sided estimators according to [4] (one-sided).
3.6.2 Field fft (field_fft)
Apply Fast Fourier Transformation (FFT) to the field data along a specified axis.
3.6.3 Dynamic mode decomposition (DMD) (field_fft)
Apply Dynamic Mode Decomposition (DMD) to the field data.
  This submodule make use of pydmd, which is not a standard requirement of pyCFS. The additional dependencies can be installed via pip using pip install pycfs[dmd].
4 Compatibility with other data formats (extras)
Library of modules to read from, convert to, and write in various formats.
4.1 Ansys Mechanical (ansys_io)
Module containing data processing utilities for reading Ansys RST files. Highlights currently include:
  • Reading Ansys result files (*.rst) into CFSMeshData and CFSResultContainer objects.
  This submodule make use of Ansys Data Processing Framework (DPF), which is not a standard requirement of pyCFS. The additional dependencies can be installed via pip using pip install pycfs[ansys].
  This submodule make use of Ansys Data Processing Framework (DPF), which is in Beta development stage. Therefore, compatibility might break rapidly.
  The methods in this submodule make use of Ansys Data Processing Framework (DPF), which requires a licensed installation of the according DPF Server!
4.2 CGNS reader (cgns_io)
Module for reading mesh files in CGNS file format, e.g. exported from ANSYS Mechanical.
  Only HDF5 based CGNS files can be read. Older CGNS versions also utilise the ADF file format, which must be converted to HDF5 beforehand. This can be done e.g. on a Linux system using cgnslib with the command: adf2hdf \$CGNS_FILE
4.3 EnSight Case Gold (ensight_io)
Module containing data processing utilities for reading EnSight Case Gold files, e.g. exported from common CFD software like openFOAM, Ansys Fluent, or StarCCM+. Highlights currently include:
  • Reading EnSight Case Gold files into numpy arrays (Mesh and Time series)
  • Conversion to CFSMeshData and CFSResultContainer objects.
  This submodule make use of VTK, which is not a standard requirement of pyCFS. The additional dependencies can be installed via pip using pip install pycfs[vtk].
4.4 Exodus file format (exodus_io)
Module for reading mesh files in Exodus file format, e.g. generated with Coreform Cubit.
4.5 MeshIO interface (meshio_io)
Load meshes via meshio into PyCFS’s CFSMeshData format.
  This submodule make use of meshio, which is not a standard requirement of pyCFS. The additional dependencies can be installed via pip using pip install pycfs[meshio].
4.6 Polytec PSV export data (psv_io)
Module containing data processing utilities for reading universal files (UFF) exported from PSV Software. Highlights currently include:
  • Reading universal files (UFF) into dict format compatible to sdypy-EMA444github.com/sdypy/sdypy-EMA.
  • Trilaterate scanning head position from export data.
  • Combine three 1D measurements into 3D data.
  • Conversion to and from CFSMeshData and CFSResultContainer objects into sdypy-EMA compatible dict format.
4.7 NiHu Matlab export data (nihu_io)
Module containing data processing utilities for reading NiHu [2] Matlab structures. Highlights currently include:
  • Conversion of NiHu mesh Matlab structure into CFSMeshData object.
4.8 STL reader (stl_io)
Function for reading mesh data from STL files.
5 Applications
Recently, pyCFS-data was used in hybrid experimental-numerical data processing routines to estimate sound power radiation of shell and plate structures including a Gaussian Process Regression [15]. Furthermore, pyCFS-data is heavily used in the pre-processing of aeroacoustic source terms of the aeroacoustic wave equation based on Pierce operator [7, 11, 8] and the compressible perturbed convective wave equation [12, 6]. It is a Python alternative of the develope C++ alternative openCFS-data [10]
Acknowledge
We would like to acknowledge the authors of pyCFS and openCFS.
Funding information
A.W. and S. S. received funding from the COMET project ECHODA ("Energy Efficient Cooling and Heating of Domestic Appliances") under FFG project No. FO999913905. ECHODA is funded within the framework of "COMET - Competence Centers for Excellent Technologies" by BMIMI, BMAW, the province of Styria, and SFG. The COMET program is managed by FFG. S. S. received funding from the FFG under project No. FO999913972.